# MGPATH: Vision-Language Model with Multi-Granular Prompt Learning for Few-Shot WSI Classification

### v3

@@ Abstract
Whole slide pathology image classification presents challenges due to gigapixel image sizes and limited annotation labels, hindering model generalization. This paper introduces a prompt learning method to adapt large vision-language models for few-shot pathology classification. We first extend the Prov-GigaPath vision foundation model, pre-trained on 1.3 billion pathology image tiles, into a vision-language model by adding adaptors and aligning it with medical text encoders via contrastive learning on 923K image-text pairs. The model is then used to extract visual features and text embeddings from few-shot annotations and fine-tunes with learnable prompt embeddings. Unlike prior methods that combine prompts with frozen features using prefix embeddings or self-attention, we propose multi-granular attention that compares interactions between learnable prompts with individual image patches and groups of them. This approach improves the model’s ability to capture both fine-grained details and broader context, enhancing its recognition of complex patterns across sub-regions. To further improve accuracy, we leverage (unbalanced) optimal transport-based visual-text distance to secure model robustness by mitigating perturbations that might occur during the data augmentation process. Empirical experiments on lung, kidney, and breast pathology modalities validate the effectiveness of our approach; thereby, we surpass several of the latest competitors and consistently improve performance across diverse architectures, including CLIP, PLIP, and Prov-GigaPath integrated PLIP. We release our implementations and pre-trained models at this MGPATH.
@@
Anne-Christin Hauschild 1,2
@@ 2 University Medical Center Göttingen, Germany
3 Max Planck Institute for Multidisciplinary Sciences, Germany
@@ 8 German Research Center for Artificial Intelligence (DFKI), Germany
Equal second contribution
@@ 1 Introduction
Whole slide imaging (WSI) Niazi et al. (2019) has become essential in modern pathology for capturing high-resolution digital representations of entire tissue samples, enabling easier digital storage, sharing, and remote analysis Pantanowitz et al. (2011). Unlike conventional methods that depend on examining slides under a microscope, WSI provides faster, detailed structural and cellular insights essential for disease diagnosis across multiple tissue layers, which is particularly valuable in cancer screening Barker et al. (2016); Cheng et al. (2021). Nevertheless, WSIs are massive images, often containing billions of pixels Farahani et al. (2015); Song et al. (2023), making detailed annotations and analysis difficult and expensive. To tackle these challenges, machine learning techniques incorporating few-shot and weakly supervised learning have been developed Madabhushi and Lee (2016); Li et al. (2023); Lin et al. (2023); Ryu et al. (2023); Shi et al. (2024). Among these, multiple instance learning (MIL) and vision-language models (VLMs) have gained particular attention for their ability to effectively manage limited annotations and interpret complex whole-slide pathology images.
@@
Figure 1: Unlike previous methods that add prompts at prefix positions or patch-level attention - disrupting structural correlations - our MGPath framework integrates prompts at both regional and individual patch levels (multi-granular attention).
@@
In MIL Ilse et al. (2018); Xu et al. (2019a); Li et al. (2023); Lin et al. (2023); Tang et al. (2023); Shi et al. (2023), each WSI is first divided into smaller patches or instances. These instances are extracted feature embeddings using pre-trained vision encoders before being grouped into a "bag", i.e., a whole slide-level representation for the entire WSI. The MIL model mainly focuses on learning ensemble functions to identify patterns in specific patches, contributing to the overall label prediction for each bag (e.g., cancerous or non-cancerous), hence reducing the need for detailed annotations. Nonetheless, these methods often struggle to select relevant patches due to complex correlations and tissue variability Gadermayr and Tschuchnig (2024); Qu et al. (2024b). To overcome those obstacles, VLMs Lu et al. (2023); Huang et al. (2023); Ikezogwo et al. (2024); Shi et al. (2024) have emerged as a promising solution, combining slide-level visual features with textual descriptions to enrich contextual understanding and support predictions in sparse data scenarios with approaches such as zero-shot learning Xu et al. (2024); Ahmed et al. (2024). Specifically, VLMs incorporate multi-scale images Shi et al. (2024); Han et al. (2024), permitting the extraction of global and local WSI features at different resolutions. To adapt the pre-trained vision-language model efficiently, prompt learning Zhou et al. (2022b); Gao et al. (2024) is employed where learnable prompts are treated as part of the input text to guide the model, and contextual prompts Li and Liang (2021); Yao et al. (2024) are integrated into feature embeddings using a self-attention mechanism Vaswani (2017). Despite their strong classification performance across diverse tasks, these approaches still encounter certain limitations.
@@
First, (i) adapting prompt learning with frozen visual features often neglects the hierarchical relationships among learnable prompts and the visual features they interact with - specifically, the multi-granular attention between prompts to individual patches and groups of patches. This limitation lessens the model’s ability to capture interdependence across distinct scales — from fine-grained local features to broader contextual information, leading to less accurate comprehension of complex patterns in pathology images. Second, (ii) many VLMs rely on the CLIP architecture Radford et al. (2021), which was not explicitly pre-trained on pathology images, thereby limiting its adaptability in few-shot settings, especially when the architecture is primarily frozen and prompt learning is applied. While there exist recent works that have incorporated PLIP Huang et al. (2023), a model pre-trained on 200k pathology image-text pairs curated from Twitter and showed significant improvements, an open question remains whether scaling pre-training to millions or billions of pathology-specific samples could further boost performance. Lastly, (iii) most VLM models for whole-slide pathology rely on cosine similarity to align visual and textual features. This metric, however, can struggle with multiple text descriptions for sub-regions Chen et al. (2023) and with augmented data perturbations Nguyen et al. (2024b), as it lacks the precision to capture fine-grained alignments between varied image-text pairs.
In this work, we present MGPath, a novel VLM method developed to address the challenges in whole-slide pathology classification. Our approach begins by adapting Prov-GigaPath Xu et al. (2024) - one of the largest pre-trained vision models trained on 1.3 billion pathology image patches - into a vision-language model. We accomplish this through contrastive learning with a pre-trained text encoder from the PLIP model Huang et al. (2023), which was trained on approximately 200K pathology image-text pairs. To strengthen this alignment, we collected an additional 923K image-text pairs from ARCH Gamper and Rajpoot (2021), PatchGastricADC22 Tsuneki and Kanavati (2022) and Quilt-1M Ikezogwo et al. (2024) and trained adaptor-based cross-alignment Gao et al. (2024); Cao et al. (2024) between Prov-GigaPath’s visual encoder and PLIP’s text encoder. Crucially, only lightweight adaptors are updated, making this process highly parameter-efficient. To the best of our knowledge, MGPath is the first parameter-efficient vision-language model trained for pathology at this data scale — utilizing 923K image-text pairs compared to the 200K in PLIP, and further benefiting from Prov-GigaPath’s 1.3 billion sample pre-training.
Next, we leverage these pre-trained models for few-shot WSI tasks by introducing multi-granular prompt learning. First, visual embeddings and descriptive text prompts are generated for image patches at different resolutions using large language models, which have been shown to improve performance Han et al. (2024); Shi et al. (2024); Qu et al. (2024a). Unlike prior methods that concatenate or use basic attention on individual patches Li and Liang (2021); Zhou et al. (2022b); Yao et al. (2024); Shi et al. (2024), our attention integrates learnable prompts with frozen visual features at both fine- and coarse-grained perspectives (Figure 1). We represent image patches from each WSI as a spatial graph, using bounding box coordinates to enable region-level aggregation through message passing along local connections. This spatial structure is encoded as tokens within the Key-Value matrices, which interact with Query matrices derived from prompt embeddings. By directing attention from Query to Key-Value matrices across both patch and region levels, our approach effectively captures hierarchical information, enriching feature representation and selectively emphasizing features across diverse tissue areas.
Finally, to measure the distance between prompt-fused visual embedding and multiple text prompts, we resort to the optimal transport (OT) method Nguyen et al. (2021); Pham et al. (2020); Séjourné et al. (2023); Chen et al. (2023); Dong et al. (2023); Nguyen et al. (2024b); Zhan et al. (2021), providing flexibility in aligning heterogeneous data distributions. This property is beneficial in few-shot WSI classification when it can (i) handle data augmentation with noise, as OT can adapt to perturbations without losing meaningful structural relationships, and (ii) capture imbalances masses between two modality embeddings when text prompts only describe sub-regions in WSI samples. Through extensive evaluations of three datasets with various architectures (CLIP-ResNet50, CLIP-ViTB16, PLIP, and (Prov-GigaPath)-integrated PLIP), we observe that MGPath demonstrate consistent improvements over several state-of-the-art MIL and VLM in literature ($14$ competitors). As an example, MGPath with (Prov-GigaPath)-PLIP variant outperforms MSCPT Han et al. (2024) by $5\%$ in F1 and $8\%$ in AUC on the TCGA-BRCA dataset. Additionally, it also surpasses two state-of-the-art VLMs models, CONCH Lu et al. (2024) and QUILT Ikezogwo et al. (2024), by approximately 6% in accuracy on TCGA-BRCA.
@@ 2.1 Large-scale Pre-trained Models for Pathology
Recent advancements in large-scale pre-trained models for pathology can be broadly classified into two categories. Vision models, such as Virchow Ikezogwo et al. (2024), Hibou Nechaev et al. (2024), UNI Chen et al. (2024), and Prov-GigaPath Xu et al. (2024) leverage massive pathology image datasets to learn robust visual representations. Among these, Prov-GigaPath stands out as the largest model, trained on 1.3 billion pathology image patches, and excels in resolving complex tissue patterns at high resolution. On the other hand, vision-language models (VLMs) like PLIP Huang et al. (2023) (trained 200K image-text pairs), CONCH Lu et al. (2024) (1.17M), or QUILTNETIkezogwo et al. (2024) (1M), integrate visual and textual information to enhance contextual understanding and improve pathology slide interpretation. In contrast, our MGPath combines the strengths of both approaches by using a parameter-efficient adaptor to link Prov-GigaPath (the largest pre-trained vision encoder) with a text encoder from VLMs like PLIP or CONCH, leveraging both rich visual features and semantic textual embeddings. Although we use the PLIP text encoder in our experiments due to its common use in baselines, the method can be extended to other larger pre-trained text models.
@@ 2.2 Few-shot learning in WSI
MIL treats a WSI as a bag of patches and aggregates these instances into a bag of features, with early methods using non-parametric techniques like mean or max pooling. However, since disease-related patches are rare, these methods can overwhelm useful information with irrelevant data. To address this, attention-based methods, graph neural Networks (GNNs), and Transformer-based methods have been introduced Lu et al. (2021); Chen et al. (2021); Ilse et al. (2018); Li et al. (2021); Shao et al. (2021); Zheng et al. (2022). In contrast, VLMs have gained popularity through contrastive learning, aligning image-text pairs to enhance performance on a variety of tasks. While collecting large-scale pathology image-text pairs remains challenging, models like MI-Zero, PLIP, and CONCH have been trained on hundreds of thousands to over a million pathology image-text pairs Lu et al. (2023); Huang et al. (2023); Lu et al. (2024). Some approaches also integrate multi-magnification images and multi-scale text to mimic pathologists’ diagnostic processes, especially for detecting subtle abnormalities Shi et al. (2024); Han et al. (2024). Our MGPath extends on the VLMs strategy by further amplifying the benefits of using a large pre-trained pathology VLM model and introducing a new parameter-efficient multi-granular prompt learning to adapt these models to few-shot settings.
@@ 2.3 Prompt Learning for Vision-Language Adaptations
Prompt tuning is proposed to transfer large pre-trained model task-specific downstream tasks and has shown strong results in multimodal models like CLIP. Rather than design a heuristic template, several methods like CoOp Zhou et al. (2022b), CoCoOp Zhou et al. (2022a), or MaPLe Khattak et al. (2023) among others Rao et al. (2022); Shu et al. (2022) have allowed models to determine optimal prompts from multiple perspectives, such as domain generalization Ge et al. (2023); Yao et al. (2024), knowledge prototype Zhang et al. (2022b); Li et al. (2024), or diversity Lu et al. (2022); Shu et al. (2022). However, these approaches focus on natural images and do not address the unique challenges of whole-slide pathology images, which require multi-scale and structural contextual information. While a few current methods typically integrate prompts with frozen visual features via self-attention Shi et al. (2024); Qu et al. (2024a), these approaches might struggle with the complex relationships in WSIs. Our solution introduces multi-granular prompt learning, bridging attention on both individual image patches and spatial groups to better align with the hierarchical structure of WSI data.
@@ 3 Methods
Figure 2 provides an overview of the key steps in our method. Before diving into the details of each section, we first introduce our PLIP model enhanced by Prov-GigaPath through the use of adaptors.
Figure 2: The pipeline of the proposed MGPath method. Low- and high-resolution image patches are processed with large language models to generate visual contextual descriptions (Section 3.2). Visual prompts are integrated with frozen features through multi-granular attention at both patch and group-of-patch levels 3.3. The final output is obtained by aligning visual and text embeddings using optimal transport (Section 3.4).

### v5

@@ Abstract
Whole slide pathology image classification presents challenges due to gigapixel image sizes and limited annotation labels, hindering model generalization. This paper introduces a prompt learning method to adapt large vision-language models for few-shot pathology classification. We first extend the Prov-GigaPath vision foundation model, pre-trained on 1.3 billion pathology image tiles, into a vision-language model by adding adaptors and aligning it with medical text encoders via contrastive learning on 923K image-text pairs. The model is then used to extract visual features and text embeddings from few-shot annotations and fine-tunes with learnable prompt embeddings. Unlike prior methods that combine prompts with frozen features using prefix embeddings or self-attention, we propose multi-granular attention that compares interactions between learnable prompts with individual image patches and groups of them. This approach improves the model’s ability to capture both fine-grained details and broader context, enhancing its recognition of complex patterns across sub-regions. To further improve accuracy, we leverage (unbalanced) optimal transport-based visual-text distance to secure model robustness by mitigating perturbations that might occur during the data augmentation process. Empirical experiments on lung, kidney, and breast pathology modalities validate the effectiveness of our approach; thereby, we surpass several of the latest competitors and consistently improve performance across diverse architectures, including CLIP, PLIP, and Prov-GigaPath integrated PLIP. We release our implementations and pre-trained models at this https://github.com/HauschildLab/MGPATH.
@@
Anne-Christin Hauschild 1,2,3
@@ 2 University Medical Center Göttingen, Germany
3 Justus Liebig University Giessen, Germany
@@ 8 German Research Center for Artificial Intelligence (DFKI), Germany
Equal second contribution. Corresponding author
@@ 1 Introduction
Whole slide imaging (WSI) (Niazi et al., 2019) has become essential in modern pathology for capturing high-resolution digital representations of entire tissue samples, enabling easier digital storage, sharing, and remote analysis (Pantanowitz et al., 2011). Unlike conventional methods that depend on examining slides under a microscope, WSI provides faster, detailed structural and cellular insights essential for disease diagnosis across multiple tissue layers, which is particularly valuable in cancer screening (Barker et al., 2016; Cheng et al., 2021). Nevertheless, WSIs are massive images, often containing billions of pixels (Farahani et al., 2015; Song et al., 2023), making detailed annotations and analysis difficult and expensive. To tackle these challenges, machine learning techniques incorporating few-shot and weakly supervised learning have been developed (Madabhushi & Lee, 2016; Li et al., 2023; Lin et al., 2023; Ryu et al., 2023; Shi et al., 2024). Among these, multiple instance learning (MIL) and vision-language models (VLMs) have gained particular attention for their ability to effectively manage limited annotations and interpret complex whole-slide pathology images.
@@
Figure 1: Unlike previous methods that add prompts at prefix positions or patch-level attention - disrupting structural correlations - our MGPath framework integrates prompts at both regional and individual patch levels (multi-granular attention). The tokens $t+1$, $t+2$, and $t+3$ represent the spatially adjacent patches within a larger tissue region (highlighted by the larger orange box), which together contribute to the model’s diagnostic decision. In contrast, $t+9$ represents a more distant or isolated patch froma different region.
@@
In MIL (Ilse et al., 2018; Xu et al., 2019a; Li et al., 2023; Lin et al., 2023; Tang et al., 2023; Shi et al., 2023), each WSI is first divided into smaller patches or instances. These instances are extracted feature embeddings using pre-trained vision encoders before being grouped into a "bag", i.e., a whole slide-level representation for the entire WSI. The MIL model mainly focuses on learning ensemble functions to identify patterns in specific patches, contributing to the overall label prediction for each bag (e.g., cancerous or non-cancerous), hence reducing the need for detailed annotations. Nonetheless, these methods often struggle to select relevant patches due to complex correlations and tissue variability (Gadermayr & Tschuchnig, 2024; Qu et al., 2024b). To overcome those obstacles, VLMs (Lu et al., 2023; Huang et al., 2023; Ikezogwo et al., 2024; Shi et al., 2024) have emerged as a promising solution, combining slide-level visual features with textual descriptions to enrich contextual understanding and support predictions in sparse data scenarios with approaches such as zero-shot learning (Xu et al., 2024; Ahmed et al., 2024). Specifically, VLMs incorporate multi-scale images (Shi et al., 2024; Han et al., 2024), permitting the extraction of global and local WSI features at different resolutions. To adapt the pre-trained vision-language model efficiently, prompt learning (Zhou et al., 2022b; Gao et al., 2024) is employed where learnable prompts are treated as part of the input text to guide the model, and contextual prompts (Li & Liang, 2021; Yao et al., 2024) are integrated into feature embeddings using a self-attention mechanism (Vaswani, 2017). Despite their strong classification performance across diverse tasks, these approaches still encounter certain limitations.
@@
First, (i) adapting prompt learning with frozen visual features often neglects the hierarchical relationships among learnable prompts and the visual features they interact with - specifically, the multi-granular attention between prompts to individual patches and groups of patches. This limitation lessens the model’s ability to capture interdependence across distinct scales — from fine-grained local features to broader contextual information, leading to less accurate comprehension of complex patterns in pathology images. Second, (ii) many VLMs rely on the CLIP architecture (Radford et al., 2021), which was not explicitly pre-trained on pathology images, thereby limiting its adaptability in few-shot settings, especially when the architecture is primarily frozen and prompt learning is applied. While there exist recent works that have incorporated PLIP (Huang et al., 2023), a model pre-trained on 200k pathology image-text pairs curated from Twitter and showed significant improvements, an open question remains whether scaling pre-training to millions or billions of pathology-specific samples could further boost performance. Lastly, (iii) most VLM models for whole-slide pathology rely on cosine similarity to align visual and textual features. This metric, however, can struggle with multiple text descriptions for sub-regions (Chen et al., 2023) and with augmented data perturbations (Nguyen et al., 2024b), as it lacks the precision to capture fine-grained alignments between varied image-text pairs.
In this work, we present MGPATH, a novel vision-language model (VLM) designed to address the challenges of whole-slide pathology classification. Unlike existing approaches, our method integrates large-scale pathology pre-training with multi-granular prompt learning to capture both fine- and coarse-grained tissue characteristics. To enhance alignment between visual and textual representations, we adopt a parameter-efficient strategy that scales to millions of image–text pairs without requiring full model retraining. Next, we incorporate optimal transport (OT) method to measure the distance between prompt-fused visual embedding and multiple text prompts, providing flexibility in aligning heterogeneous data distributions.
Extensive evaluations on three benchmark datasets and across diverse backbones demonstrate that MGPATH consistently outperforms both multiple instance learning (MIL) and recent VLM approaches, achieving significant gains in F1, AUC, and accuracy over state-of-the-art baselines. Our contributions are summarized as follows:
  • We propose MGPATH, a parameter-efficient vision–language model for whole-slide pathology that leverages large-scale pathology pretraining with lightweight adaptor-based alignment.
  • We introduce a multi-granular prompt learning framework with hierarchical attention, enabling effective integration of fine- and coarse-grained contextual information from whole-slide images.
  • We incorporate optimal transport to align prompt-fused visual embeddings and multiple textual prompts, providing robustness to align heterogeneous data distributions in few-shot scenarios.
  • We evaluate MGPATH across three pathology benchmarks and multiple backbones, consistently outperforming state-of-the-art MIL and VLM baselines, with notable improvements in F1, AUC, and accuracy.
@@ 2.1 Large-scale Pre-trained Models for Pathology
Recent advancements in large-scale pre-trained models for pathology can be broadly classified into two categories. Vision models, such as Virchow (Ikezogwo et al., 2024), Hibou (Nechaev et al., 2024), UNI (Chen et al., 2024), and Prov-GigaPath (Xu et al., 2024) leverage massive pathology image datasets to learn robust visual representations. Among these, Prov-GigaPath stands out as the largest model, trained on 1.3 billion pathology image patches, and excels in resolving complex tissue patterns at high resolution. On the other hand, vision-language models (VLMs) like PLIP (Huang et al., 2023) (trained 200K image-text pairs), CONCH (Lu et al., 2024) (1.17M), or QUILTNET (Ikezogwo et al., 2024) (1M), integrate visual and textual information to enhance contextual understanding and improve pathology slide interpretation. In contrast, our MGPATH combines the strengths of both approaches by using a parameter-efficient adaptor to link Prov-GigaPath (the largest pre-trained vision encoder) with a text encoder from VLMs like PLIP or CONCH, leveraging both rich visual features and semantic textual embeddings. Although we use the PLIP text encoder in our experiments due to its common use in baselines, the method can be extended to other larger pre-trained text models.
@@ 2.2 Few-shot learning in WSI
MIL treats a WSI as a bag of patches and aggregates these instances into a bag of features, with early methods using non-parametric techniques like mean or max pooling. However, since disease-related patches are rare, these methods can overwhelm useful information with irrelevant data. To address this, attention-based methods, graph neural Networks (GNNs), and Transformer-based methods have been introduced (Lu et al., 2021; Chen et al., 2021; Ilse et al., 2018; Li et al., 2021; Shao et al., 2021; Zheng et al., 2022). In contrast, VLMs have gained popularity through contrastive learning, aligning image-text pairs to enhance performance on a variety of tasks. While collecting large-scale pathology image-text pairs remains challenging, models like MI-Zero, PLIP, and CONCH have been trained on hundreds of thousands to over a million pathology image-text pairs (Lu et al., 2023; Huang et al., 2023; Lu et al., 2024). Some approaches also integrate multi-magnification images and multi-scale text to mimic pathologists’ diagnostic processes, especially for detecting subtle abnormalities (Shi et al., 2024; Han et al., 2024). Our MGPATH extends on the VLMs strategy by further amplifying the benefits of using a large pre-trained pathology VLM model and introducing a new parameter-efficient multi-granular prompt learning to adapt these models to few-shot settings.
@@ 2.3 Prompt Learning for Vision-Language Adaptations
Prompt tuning is proposed to transfer large pre-trained model task-specific downstream tasks and has shown strong results in multimodal models like CLIP. Rather than design a heuristic template, several methods like CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), or MaPLe (Khattak et al., 2023) among others (Rao et al., 2022; Shu et al., 2022) have allowed models to determine optimal prompts from multiple perspectives, such as domain generalization (Ge et al., 2023; Yao et al., 2024), knowledge prototype (Zhang et al., 2022b; Li et al., 2024), or diversity (Lu et al., 2022; Shu et al., 2022). However, these approaches focus on natural images and do not address the unique challenges of whole-slide pathology images, which require multi-scale and structural contextual information. While a few current methods typically integrate prompts with frozen visual features via self-attention (Shi et al., 2024; Qu et al., 2024a), these approaches might struggle with the complex relationships in WSIs. Our solution introduces multi-granular prompt learning, bridging attention on both individual image patches and spatial groups to better align with the hierarchical structure of WSI data.
@@ 3 Methods
This section presents MGPATH, a parameter-efficient vision–language framework for whole-slide pathology classification. We bridge the Prov-GigaPath visual encoder (Xu et al., 2024) - pretrained on 1.3B pathology patches—and the PLIP text encoder (Huang et al., 2023) - trained on $\sim$200K image–text pairs—using lightweight adaptor modules optimized with a contrastive objective. To strengthen vision–text alignment without updating either backbone, we collect an additional 923K pathology image–text pairs from ARCH (Gamper & Rajpoot, 2021), PatchGastricADC22 (Tsuneki & Kanavati, 2022), and Quilt-1M (Ikezogwo et al., 2024) and train only the adaptors via cross-alignment (Gao et al., 2024; Cao et al., 2024). This yields a large-scale aligned representation while remaining highly parameter-efficient. To our knowledge, MGPATH is the first parameter-efficient VLM for pathology trained at this data scale (923K pairs) on top of a 1.3B-patch visual pretraining.
Next, we leverage these pre-trained models for few-shot WSI tasks by introducing multi-granular prompt learning. First, descriptive text prompts are generated for image patches at dual magnification (low/high) using a frozen large language model (LLM) and augment them with multiple learnable prompt tokens per magnification, capturing diverse subregional patterns (Han et al., 2024; Shi et al., 2024; Qu et al., 2024a). Then, we integrate the learnable prompts with frozen visual features at two granularities: (i) patch-level attention over all patches and (ii) region-level attention over spatial groups built from patch coordinates via message passing. As illustrated in Figure 1, unlike previous methods, which typically score tokens independently, based on their individual relevance scores to the prompts ($t+3$ and $t+9$). Our method explicitly incorporates both regional attention adjacent, semantically linked tokens ($t+1$, $t+2$, $t+3$), while preserving patch-level attention for isolated yet discriminative tokens ($t+9$). This results in richer prompt-guided slide features that better capture the underlying pathology. Concretely, we represent image patches from each WSI as a spatial graph, using bounding-box coordinates to enable region-level aggregation through message passing along local connections. This spatial structure is encoded as tokens within the Key-Value matrices, which interact with Query matrices derived from prompt embeddings. By directing attention from Query to Key-Value matrices across both patch and region levels, our approach effectively captures hierarchical information, enriching feature representation and selectively emphasizing features across diverse tissue areas. Finally, instead of cosine similarity, we align prompt-fused visual summaries to class prompts using an optimal transport (OT) distance (Nguyen et al., 2021; Pham et al., 2020; Séjourné et al., 2023; Chen et al., 2023; Dong et al., 2023; Nguyen et al., 2024b; Zhan et al., 2021), which is robust to noisy augmentations and partial text–region overlap - both common in WSIs.
Figure 2 provides an overview of the key steps in our method. We next detail the adaptor-based vision–text alignment (Sec. 3.1), the construction of multi-magnification descriptive prompts (Sec. 3.2), granularity-aware visual prompting (Sec. 3.3), and the OT-based alignment objective (Sec. 3.4).
Figure 2: The pipeline of the proposed MGPATH. Low- and high-resolution image patches are processed with large language models to generate visual contextual descriptions (Section 3.2). Visual prompts are integrated with frozen features through multi-granular attention at both patch and group-of-patch levels $3.$ The final output is obtained by aligning visual and text embeddings using optimal transport (Section 3.4).